1/2(3a+1)-1/3(a-5)=

Answer by cngriffith(27)

1/2(3a=1)-1/3(a-5)=
(3a+1)/2-(a-5)/3= Use distributive property to multiply
Now fine the common denominator of 2 and 3 which is 6.
Change each complex fraction into an equivalent one
with a denominator of 6.
3(3a+1)-2(a-5)/6 Use distributive property to multiply the numerator.
9a+3-2a+5/6 The result after multiplying.
7a+8/6 Combine like terms in the numerator.
7a/6 +4/3 Divide each term of the numerator by the 6 in the
denominator and reduce where possible.